OpenStudy (anonymous):

find the solution to the system of equations by the addition method. x-3y = -25 8x - 3y = -74

hero (hero):

Want my solution?

OpenStudy (anonymous):

yea

hero (hero):

I would rather you understand how to do it.

hero (hero):

All you have to do is simply 1. isolate y, 2. set y = y, 3. solve for x 4. solve for y

OpenStudy (anonymous):

whats a easy method? sry i just got surgery and my brain isn't working as well as i want it too. im doing some homework

hero (hero):

That IS the easy method. All you have to do is just realize it and try to understand it.

hero (hero):

If it wasn't the easy method, I wouldn't be wasting my time showing you.

OpenStudy (anonymous):

thank u i appreciate it

hero (hero):

If you just isolate y, you could actually plug that into the calculator and find the intersecting point.

hero (hero):

When you solve for y, it will have the form y = mx+b, but I usually solve these algebraically, and to check myself, I can used calc if I choose.

OpenStudy (anonymous):

i wish my school taught me how to use this calculator right. math is so frusterating

OpenStudy (anonymous):

i really appreciate ur help tho

hero (hero):

0. Original Problem: x-3y = -25 8x - 3y = -74 1. Isolate y: y = (25 + x)/3 y = (74 + 8x)/3 y = y 25 + x = 74 + 8x 25 - 74 = 8x - x -49 = 7x -49/7 = x -7 = x 6 = y

hero (hero):

Do you see where y is isolated? y = (25 + x)/3 y = (74 + 8x)/3 ^^^^^^^^ You can use your calc to input those two expressions and it will give you the intersection point.